# Latent Diffusion Inversion Requires Understanding the Latent Space

### v1

@@ Abstract
The recovery of training data from generative models (“model inversion”) has been extensively studied for diffusion models in the data domain. The encoder/decoder pair and corresponding latent codes have largely been ignored by inversion techniques applied to latent space generative models, e.g., Latent Diffusion models (LDMs). In this work we describe two key findings: (1) The diffusion model exhibits non-uniform memorization across latent codes, tending to overfit samples located in high-distortion regions of the decoder pullback metric. (2) Even within a single latent code, different dimensions contribute unequally to memorization. We introduce a principled method to rank latent dimensions by their per-dimensional contribution to the decoder pullback metric, identifying those most responsible for memorization. Empirically, removing less-memorizing dimensions when computing attack statistics for score-based membership inference attacker significantly improves performance, with average AUROC gains of 2.7% and substantial increases in TPR@1%FPR (6.42%) across diverse datasets including CIFAR-10, CelebA, ImageNet-1K, Pokémon, MS-COCO, and Flickr. This indicates stronger confidence in identifying members under extremely low false-positive tolerance. Our results highlight the overlooked influence of the auto-encoder geometry on LDM memorization and provide a new perspective for analyzing privacy risks in diffusion-based generative models.
@@ 1 Introduction
Model inversion is the task of recovering training samples from a trained generative model. The generative models ostensibly should be fitting an underlying data distribution, so finding the individual training points implies an overfitting, and moreover may be informative of pathological structures in fitted models. The inversion of diffusion models has been extensively explored in prior work [zhu2016generative, creswell2018inverting, carlini2023extracting, somepalli2023diffusion, somepalli2023understanding, gu2023memorization]. However, for Latent Diffusion Models (LDMs) [rombach2022high], existing research at onces (1) predominantly focuses on inverting the diffusion process itself, treating the latent space as a fixed substrate and largely overlooking the role of the associated variational auto-encoder (VAE), and (2) has significantly reduced inversion performance [rao2025score] relative to data-domain diffusion processes.
@@
A crucial precursor to model inversion is the membership inference attack (MIA), where attack methods attempt to distinguish training examples (members) from non-training examples (non-members). This measures how well we can identify training-points without considering the relatively unrelated search problem over the space of images. Many effective MIA methods have been proposed for diffusion models [carlini2023extracting, zhai2024membership, kong2023efficient, matsumoto2023membership, carlini2023extracting, fu2023probabilistic]. A model with stronger training-set memorization (overfit) typically shows higher MIA vulnerability; thus, the sensitivity to MIA serves as a practical metric for evaluating memorization.
@@
Prior work [rao2025score] shows that varying the latent-space regularization leads to substantially different degrees of vulnerability to membership inference, indicating that the structure of the latent space influences model memorization. In this paper we take this a step further, demonstrating that geometric properties of the decoder implicate upon membership inference performance. Based on statistics of the decoder geometry, we construct a filtering procedure that improves membership inference performance across all tested methods; this demonstrates that these local statistics are connected with overfit phenomena, and not simply biases or idiosyncrasies in particular model inversion techniques.
@@
To characterize geometric properties of the decoder, we employ the pullback metric from Riemannian geometry [arvanitidis2017latent, chadebec2022geometric, arvanitidis2020geometrically]. This is metric tensor associated with the decoder mapping (the “pulling back” of the data-space metric to the latent space), and its determinant measures the local volume change induced by the decoder [arvanitidis2017latent], which we call the distortion. Empirically, we find that latent diffusion models tend to memorize samples located in regions of high decoder distortion more strongly than those in low-distortion regions (diagrammed in Fig. 1). In contrast, if the latent space exhibits nearly uniform distortion, memorization becomes correspondingly uniform. The strong correlation between decoder’s local distortion and memorization indicates that the choice of encoder/decoder architecture plays an important role in score field learning. Experiments to support these claims are presented in Section 4.2.
@@
Previous work observed that data domain diffusion denoisers preferentially reconstruct high-variance directions while oversmoothing low-variance directions [lukoianov2025locality, wang2023hidden, kadkhodaie2023generalization]. These directions correspond to eigenvectors of the data covariance matrix. Although this analogy is not rigorously transported to the latent space without significant computation, whose nonlinear encoding behavior can differ substantially point-to-point, it suggests that certain latent directions may contribute disproportionately to memorization. We improve the inversion performance by preferentially including or excluding different latent space dimensions based on the per-dimensional contribution to the local distortion induced by the decoder. In Section 3.3, we introduce a computationally efficient method to quantify these contributions, diagrammed in Fig. 1 (right). We empirically validate this idea by selectively removing the less-contributing (alternatively, less-memorizing) dimensions and evaluating the resulting membership inference attacks. Across multiple datasets, this dimensional masking consistently improves attack performance: AUROC increases by 1–4%, and TPR@1%FPR improves by 1–15%.
@@
All data splits, model checkpoints, training/fine-tuning scripts, and testing code are released on our GitHub repository. [withheld for review]
@@ 2 Related Work
Riemannian Geometry of Variational Autoencoders.
@@
A line of work studies the Riemannian geometry induced by deep generative models, and VAEs in particular, via the pullback metric of the decoder. Arvanitidis et al. show that the generator of a VAE induces a stochastic Riemannian metric on the latent space and that geodesics and distances computed under this metric lead to more semantically meaningful interpolations, improved sampling, and better clustering [arvanitidis2017latent, arvanitidis2020geometrically]. shao2018riemannian further develop algorithms for computing geodesics and parallel transport on the manifolds defined by deep generative models. Subsequent work has used this geometric viewpoint to regularize latent spaces, design geometry-aware VAEs, and perform data augmentation in small-sample regimes by exploiting the learned Riemannian structure [chadebec2022geometric, arvanitidis2020geometrically]. Different from these works, which primarily use the pullback metric to study interpolation, representation quality, and manifold geometry, our goal is different: we treat the decoder-induced pullback metric as a privacy-relevant signal. We estimate local distortion for high-dimensional VAEs used in LDMs, and we show that this distortion reveals where the latent diffusion model memorizes most strongly and which latent dimensions contribute most to membership leakage.
Latent-domain diffusion models.
@@
LDMs [rombach2022high] factor generative modeling into a VAE and a diffusion model trained in the resulting latent space, dramatically reducing computational cost while preserving visual fidelity. Subsequent work has largely followed this paradigm but focused on improving synthesis quality, efficiency, or conditioning, for example by replacing the standard VAE with alternative autoencoder families or representation autoencoders [zheng2025diffusion, shi2025latent], or by scaling latent diffusion to new modalities and tasks. In parallel, a growing literature analyzes privacy and memorization in diffusion models, proposing increasingly strong membership inference and data extraction attacks [carlini2023extracting, duan2023diffusion, kong2023efficient, fu2023probabilistic, zhai2024membership, pang2023white], and several works study frequency-domain behavior and “pass-through” phenomena of data-domain denoisers [lukoianov2025locality, wang2023hidden, kadkhodaie2023generalization, li2024unveiling, lian2025unveiling]. However, these analyses either operate in pixel space or treat the autoencoder in LDMs as a fixed preprocessor. By contrast, we explicitly attribute memorization behavior of latent-space diffusion models to the geometry of the encoder–decoder pair. Our results show that decoder-induced local distortion in the VAE latent space controls spatially non-uniform memorization and that per-dimension contributions to this distortion identify latent directions that dominate membership leakage, bringing attention to the encoder–decoder architecture as a primary factor in the inversion properties of LDMs.
Frequency analysis on data-domain diffusion models.
@@
Prior work [lukoianov2025locality, wang2023hidden] has shown that pixel-domain diffusion model denoisers often struggle with high-frequency components, which correspond to low-variance directions in the eigen-space of the data covariance. In particular, these denoisers tend to suppress high-frequency content and implicitly project reconstructions toward high-variance, low-frequency directions. However, in latent-domain diffusion models, the dynamics are fundamentally different: the diffusion process operates on nonlinear, semantically structured latent representations produced by an encoder, and the mapping back to the image domain is mediated by a decoder with nontrivial geometry. We empirically show that the local distortion of an image in latent space is not fully explained by its high-frequency energy in the pixel domain (see Table 3). This suggests that frequency-based analyses developed for data-domain diffusion models [li2024unveiling, lian2025unveiling] may not directly transfer to latent-domain models.
@@
Moreover, lian2025unveiling report that suppressing high-frequency components in the denoised images can enhance the separation between membership attack statistics on training versus non-training data. In our experiments with Stable Diffusion, we do not observe a similar improvement when applying analogous high-frequency suppression (see Appendix, Section E). Taken together, these findings indicate that latent-domain diffusion models require a different analytical perspective: rather than relying solely on frequency-based reasoning in pixel space, one must account for the nonlinear encoder–decoder geometry and its induced distortion in latent space.
@@
Figure 2: Each violin plot shows the distribution of the decoder-induced local distortion, with representative samples displayed at high-distortion (top) and low-distortion (bottom) regions. CelebA shows an almost uniform distortion distribution, while CIFAR-10 and ImageNet-1K display larger variation across samples. Refer to the appendix for additional samples.
@@ 3 Methodology
Terminology. Unless otherwise specified, the term latent space refers to the latent representation learned by the VAE component of LDM. The diffusion process operating in this latent representation is referred to as the latent space diffusion model.
@@ 3.1 Score-based Membership Inference Attacks
Membership Inference Attacks (MIA) are pre-cursor tasks to model inversion. They measure the separability (conditional on an attack methodology/model) of seen training and unseen test data given a trained diffusion model and a threat model. While not performing the full search required by model inversion, MIA demonstrates that candidate data-points can be classified as in training or out of training. For our threat model, we consider a common grey-box [duan2023diffusion, zhai2024membership, kong2023efficient, matsumoto2023membership, carlini2023extracting, fu2023probabilistic] scenario, where the attack can manipulate model inputs but should be agnostic to particular architectures or weight configurations.
@@
often averaged over several $\varepsilon$ draws [matsumoto2023membership]. SecMI further exploits diffusion structure by measuring a single–step posterior estimation error (“$t$-error”) at a chosen timestep $t$ using deterministic DDIM mappings. Let $\Phi_{\theta}$ and $\Psi_{\theta}$ denote the deterministic reverse/denoise operators; with $\tilde{x}_{t}=\Phi_{\theta}(x^{\star},t)$, SecMI’s statistic is
@@
which is thresholded (or fed to a small attack network) to decide membership [duan2023diffusion]. PIA attains similar accuracy with far fewer queries by using the model’s own t=0t{=}0 denoiser output as a proximal initialization and then computing a forward/backward consistency error built from this initialization [kong2023efficient]. For $\tilde{x}_{t}^{\star}=\sqrt{\bar{\alpha}_{t}}x^{*}+\sqrt{1-\bar{\alpha}_{t}}\hat{\varepsilon}_{\theta}(x,t=0)$,
$\text{PIA}(x^{\star},t)=\|\hat{\varepsilon}_{\theta}(x^{\star},t=0)-\hat{\varepsilon}_{\theta}(\tilde{x}_{t}^{\star},t)\|$ (1)
Finally, SimA [rao2025score] is the scaled Loss criteria at $\varepsilon=0$ (point estimate), which is the mean and mode of $\varepsilon$ distribution. This is inherently a direct score estimation.
@@
motivated by theory linking the expected denoiser to a kernel–weighted local mean of nearby training samples. Empirically, SimA matches or exceeds multi–query baselines while being simpler and faster; hence we adopt SimA to quantify the model memoriztion in most experiments [rao2025score].
@@ 3.2 VAE Decoder Pullback Metric
that transforms a latent variable $z\in\mathcal{Z}$ into a data point $x=D(z)\in\mathcal{X}$. While $\mathcal{Z}$ is usually assumed to have Euclidean structure, this assumption is also usually incorrect when viewed in terms of the decoder and $\mathcal{X}$: the nonlinear decoder distorts distances, making Euclidean geometry inapplicable. Instead, the decoder induces a Riemannian metric on $\mathcal{Z}$, known as the pullback metric [arvanitidis2017latent, arvanitidis2020geometrically, shao2018riemannian].
@@
Large singular values appear to correspond to directions that produce fine-grained, spatially detailed (high-frequency) pixel changes. Note that in VAE theory, there is little to no reason that this should be true, but our empirical evidence across datasets shows it to be a consistent trend. If one image’s $S_{K}$ is much larger, Fig. 2 shows this image is usually semantically meaningful, which we assume that image’s local neighborhood contains more high-frequency detail — edges, texture — while the other image’s neighborhood is smoother or simpler. More examples can be found in Appendix F. Nevertheless, to our knowledge, no prior work demonstrates that greater distortion implies greater semantic variation.
@@ 3.3 Quantifying Per-dimension Influence on Latent Memorization
Hutchinson Estimator: Instead of explicitly computing the diagonal entries of the pullback metric $G_{ii}(z)=\|J_{D}(z)e_{i}\|_{2}^{2}$, which requires constructing the full decoder Jacobian $J_{D}(z)$, we estimate them efficiently using the Hutchinson stochastic trace estimator. For any function $D:\mathcal{Z}\!\to\!\mathcal{X}$ with Jacobian $J_{D}(z)$, the diagonal of the Gram matrix can be expressed as
@@
This estimator takes $\mathcal{O}(n_{\mathrm{mc}}\,T_{D})$, which scales linearly in both the number of latent samples and the number of Monte Carlo probes, and independent of the explicit Jacobian size. Empirically, we found $n_{\mathrm{mc}}{=}8$ is good enough for image datasets of all resolutions. A pseudocode is provided in appendix  D.

### v3

@@ Abstract
The recovery of training data from generative models (“model inversion”) has been extensively studied for diffusion models in the data domain as a memorization/overfitting phenomenon. Latent diffusion models (LDMs), which operate on the latent codes from encoder/decoder pairs, have been robust to prior inversion methods. In this work we describe two key findings: (1) the diffusion model exhibits non-uniform memorization across latent codes, tending to overfit samples located in high-distortion regions of the decoder pullback metric; (2) even within a single latent code, memorization contributions are unequal across representation dimensions. Our proposed method to ranks latent dimensions by their contribution to the decoder pullback metric, which in turn identifies dimensions that contribute to memorization. For score-based membership inference, a sub-task of model inversion, we find that removing less-memorizing dimensions improves performance on all tested methods and datasets, with average AUROC gains of 1–4% and substantial increases in TPR@1%FPR (1–32%) across diverse datasets including CIFAR-10, CelebA, ImageNet-1K, Pokémon, MS-COCO, and Flickr. Our results highlight the overlooked influence of the auto-encoder geometry on LDM memorization and provide a new perspective for analyzing privacy risks in diffusion-based generative models.
@@ 1 Introduction
Model inversion is the task of recovering training samples from a trained generative model. Generative models should ostensibly be fitting an underlying data distribution; recovering individual training points from their model weights implies an overfitting or memorization phenomenon. Inversion methods may be informative of pathological structures and fitting behavior, both for fitted models and more broadly for the model family. In this paper we focus on Latent Diffusion Models (LDMs) [28] and pathological overfitting behaviors resulting in their invertibility.
@@
A crucial precursor to model inversion is the membership inference attack (MIA), where attack methods attempt to distinguish training examples (members) from non-training examples (non-members). This measures how well we can identify training-points without considering the relatively unrelated search problem over the space of images. Many effective MIA methods have been proposed for diffusion models [5, 37, 17, 23, 11]. A model with stronger training-set memorization (overfit) typically shows higher MIA vulnerability; thus, the sensitivity to MIA serves as a practical metric for evaluating memorization.
The inversion of diffusion models has been extensively explored in prior work [39, 8, 5, 33, 34, 14]. However, for Latent Diffusion Models (LDMs) [28], existing research at once (1) predominantly focuses on inverting the diffusion process itself, treating the latent space as a fixed substrate and largely overlooking the role of the associated variational auto-encoder (VAE), and (2) has significantly reduced inversion performance [27] relative to data-domain diffusion processes.
@@
Prior work [27] shows that varying the latent-space regularization leads to substantially different degrees of vulnerability to membership inference, indicating that the structure of the latent space influences model memorization. In this paper we take this a step further, demonstrating that geometric properties of the decoder implicate upon membership inference performance. Based on statistics of the decoder geometry, we construct a filtering procedure that improves membership inference performance across all tested methods; this demonstrates that these local statistics are connected with overfit phenomena, and not simply biases or idiosyncrasies in particular model inversion techniques.
@@
To characterize geometric properties of the decoder, we employ the pullback metric from Riemannian geometry [2, 6, 3]. This is metric tensor associated with the decoder mapping (the “pulling back” of the data-space metric to the latent space), and its determinant measures the local volume change induced by the decoder [2], which we call the distortion. Empirically, we find that latent diffusion models tend to memorize samples located in regions of high decoder distortion more strongly than those in low-distortion regions (diagrammed in Fig. 1). In contrast, if the latent space exhibits nearly uniform distortion, memorization becomes correspondingly uniform. The strong correlation between decoder’s local distortion and memorization indicates that the choice of encoder/decoder architecture plays an important role in score field learning. Experiments to support these claims are presented in Section 4.2.
@@
Previous work observed that data domain diffusion denoisers preferentially reconstruct high-variance directions while oversmoothing low-variance directions [22, 35, 16]. These directions correspond to eigenvectors of the data covariance matrix. Although this analogy is not rigorously transported to the latent space without significant computation, whose nonlinear encoding behavior can differ substantially point-to-point, it suggests that certain latent directions may contribute disproportionately to memorization. We improve the inversion performance by preferentially including or excluding different latent space dimensions based on the per-dimensional contribution to the local distortion induced by the decoder. In Section 3.3, we introduce a computationally efficient method to quantify these contributions, diagrammed in Fig. 1 (right). We empirically validate this idea by selectively removing the less-contributing (alternatively, less-memorizing) dimensions and evaluating the resulting membership inference attacks. Across multiple datasets, this dimensional masking consistently improves attack performance of all attack methods: AUROC increases by 1–4%, and TPR@1%FPR improves by 1–32%.
@@
All data splits, model checkpoints, training/fine-tuning scripts, and testing code are released on our GitHub repository https://github.com/mx-ethan-rao/VAE2Diffusion.git.
@@ 2 Related Work
Riemannian Geometry of Variational Autoencoders:
@@
A line of work studies the Riemannian geometry induced by deep generative models, and VAEs in particular, via the pullback metric of the decoder. Arvanitidis et al. [2] show that the generator of a VAE induces a stochastic Riemannian metric on the latent space and that geodesics and distances computed under this metric lead to more semantically meaningful interpolations, improved sampling, and better clustering [2, 3]. Shao et al. [31] further develop algorithms for computing geodesics and parallel transport on the manifolds defined by deep generative models. Subsequent work has used this geometric viewpoint to regularize latent spaces, design geometry-aware VAEs, and perform data augmentation in small-sample regimes by exploiting the learned Riemannian structure [6, 3]. These works primarily use the pullback metric to study interpolation, representation quality, and manifold geometry, while instead we study the decoder-induced pullback metric in terms of its privacy-relevant changes to the original space. We estimate local distortion for high-dimensional VAEs used in LDMs, and we show that this distortion reveals where the latent diffusion model memorizes most strongly and which latent dimensions contribute most to membership leakage.
Latent-domain diffusion models:
@@
LDMs [28] factorize generative modeling into a VAE and a diffusion model trained in the resulting latent space, dramatically reducing computational cost while preserving visual fidelity. Subsequent work has largely followed this paradigm but focused on improving synthesis quality, efficiency, or conditioning, e.g., by replacing the standard VAE with alternative autoencoder families [38, 32], or by scaling latent diffusion to new modalities and tasks. In parallel, a growing literature analyzes privacy and memorization in diffusion models, proposing increasingly strong membership inference and data extraction attacks [5, 9, 17, 11, 37, 25]; several of these membership inference studies show “pass-through” phenomena of data-domain denoisers [16, 18, 19]. However, these analyses either operate in pixel space or treat the autoencoder in LDMs as a fixed pre-processor. In contrast, we explicitly attribute memorization behavior of latent-space diffusion models to the geometry of the encoder/decoder pair. Our results show that decoder-induced local distortion in the VAE latent space controls spatially non-uniform memorization and that per-dimension contributions to this distortion identify latent directions that dominate membership leakage, bringing attention to the encoder/decoder architecture as a primary factor in the inversion properties of LDMs.
@@
Prior work [22, 35] has shown that pixel-domain diffusion model denoisers often struggle with high-frequency components, which correspond to low-variance directions in the eigen-space of the data covariance. In particular, these denoisers tend to suppress high-frequency content and implicitly project reconstructions toward high-variance, low-frequency directions. However, in latent-domain diffusion models the diffusion process operates on nonlinear, semantically structured latent representations produced by an encoder, and the mapping back to the image domain is mediated by a decoder with nontrivial geometry. We empirically show that the local distortion of an image in latent space is not fully explained by its high-frequency energy in the pixel domain (see Table 3). This suggests that frequency-based analyses developed for data-domain diffusion models [18, 19] may not directly transfer to latent-domain models, though there is a analogous non-linear effect.
@@
Moreover, Lian et al. [19] report that suppressing high-frequency components in the denoised images can enhance the separation between membership attack statistics on training versus non-training data. In our experiments with Stable Diffusion, we do not observe a similar improvement when applying analogous high-frequency suppression (see Appendix, Section F). Taken together, these findings indicate that latent-domain diffusion models require a different analytical perspective: rather than relying solely on frequency-based reasoning in pixel space, one must account for the nonlinear encoder–decoder geometry and its induced distortion in latent space.
@@
Figure 2: Each violin plot shows the distribution of the decoder-induced local distortion, with representative samples displayed at high-distortion (top) and low-distortion (bottom) regions. CelebA shows an almost uniform distortion distribution, while CIFAR-10 and ImageNet-1K display larger variation across samples. Refer to the appendix G for additional samples.
@@ 3 Methodology
Terminology: Unless otherwise specified, the term latent space refers to the latent representation learned by the VAE component of LDM. The diffusion process operating in this latent representation is referred to as the latent space diffusion model.
@@ 3.1 Score-based Membership Inference Attacks
Membership Inference Attacks (MIA) are pre-cursor tasks to model inversion. They measure the separability (conditional on an attack methodology/model) of seen training and unseen test data given a trained diffusion model and a threat model. While not performing the full search required by model inversion, MIA demonstrates that candidate data-points can be classified as in training or out of training. For our threat model, we consider a common grey-box [9, 37, 17, 23, 5, 11] scenario, where the attack can manipulate model inputs but should be agnostic to particular architectures or weight configurations.
@@
often averaged over several $\varepsilon$ draws [23]. SecMI further exploits diffusion structure by measuring a single–step posterior estimation error (“$t$-error”) at a chosen timestep $t$ using deterministic DDIM mappings. Let $\Phi_{\theta}$ and $\Psi_{\theta}$ denote the deterministic reverse/denoise operators; with $\tilde{x}_{t}=\Phi_{\theta}(x^{\star},t)$, SecMI’s statistic is
@@
which is thresholded (or fed to a small attack network) to decide membership [9]. PIA attains similar accuracy with far fewer queries by using the model’s own t=0t{=}0 denoiser output as a proximal initialization and then computing a forward/backward consistency error built from this initialization [17]. For $\tilde{x}_{t}^{\star}=\sqrt{\bar{\alpha}_{t}}x^{*}+\sqrt{1-\bar{\alpha}_{t}}\hat{\varepsilon}_{\theta}(x,t=0)$,
$\text{PIA}(x^{\star},t)=\|\hat{\varepsilon}_{\theta}(x^{\star},t=0)-\hat{\varepsilon}_{\theta}(\tilde{x}_{t}^{\star},t)\|$
Finally, SimA [27] is the scaled Loss criteria at $\varepsilon=0$ (point estimate), which is the mean and mode of $\varepsilon$ distribution. This is inherently a direct score estimation.
@@
motivated by theory linking the expected denoiser to a kernel–weighted local mean of nearby training samples. Empirically, SimA matches or exceeds multi–query baselines while being simpler and faster; hence we adopt SimA to quantify the model memoriztion in most experiments [27].
@@ 3.2 VAE Decoder Pullback Metric
that transforms a latent variable $z\in\mathcal{Z}$ into a data point $x=D(z)\in\mathcal{X}$. While $\mathcal{Z}$ is usually assumed to have Euclidean structure, this assumption is also usually incorrect when viewed in terms of the decoder and $\mathcal{X}$: the nonlinear decoder distorts distances, making Euclidean geometry inapplicable. Instead, we need to choose a more relevant metric for the space; one promising candidate from the literature is the decoder induced Riemannian metric on $\mathcal{Z}$, known as the pullback metric [2, 3, 31].
@@
Large singular values appear to correspond to directions that produce fine-grained, spatially detailed (high-frequency) pixel changes. If one image’s $S_{K}$ is much larger, Fig. 2 shows this image is usually semantically meaningful, which we assume that image’s local neighborhood contains more high-frequency detail (edges or textures) while the other image’s neighborhood is smoother or simpler. More examples and additional exposition on this phenonmenon in VAE can be found in Appendix G.
@@ 3.3 Quantifying Per-dimension Influence on Latent Memorization
Hutchinson Estimator: Instead of explicitly computing the diagonal entries of the pullback metric $G_{ii}(z)=\|J_{D}(z)e_{i}\|_{2}^{2}$, which requires constructing the full decoder Jacobian $J_{D}(z)$, we estimate them efficiently using the Hutchinson stochastic trace estimator [13]. For any function $D:\mathcal{Z}\!\to\!\mathcal{X}$ with Jacobian $J_{D}(z)$, the diagonal of the Gram matrix can be expressed as
@@
This estimator takes $\mathcal{O}(n_{\mathrm{mc}}\,T_{D})$, where $T_{D}$ denotes the computational cost of a single evaluation of the decoder $D$. It scales linearly in the number of Monte Carlo probes, and is independent of the explicit Jacobian size. Empirically, we found $n_{\mathrm{mc}}{=}8$ is good enough for image datasets of all resolutions. A pseudocode is provided in appendix  D.